OpenStudy (fernandez555):

***Reward+Fan**** Please check my answer, just tell me if im right. Given the trinomial 3x^2 - 6x + 5, what is the value of the discriminant? A. -66 B. -24 C. 8 D. 96 I put D. 96

OpenStudy (ajprincess):

It isn't right.:(. discriminant=b^2-4ac here b=-6 a=3 c=5

OpenStudy (fernandez555):

@ajprincess damn it. so, B^2 = -12 then, -12 - 4(5*3) correct?

OpenStudy (ajprincess):

b=-6 b^2=(-6)*(-6)=36 b^2-4ac=36-4(3*5)

OpenStudy (fernandez555):

@ajprincess ahh careless mistake, multiplied by 2 instead of squared. so B then, -24 : )

OpenStudy (ajprincess):

yup:)